three numbers are in arithmetic progresion & sum of their
squares is 140. find the numbers?

Answer by Edwin McCravy(20056)

There is not enough information.  Here are 8 different solutions.
These are only the integer solutions.  There are infinitely many
irrational solutions.  Look carefully at what was given and see
what is omitted.

{-2,-6,-10} and (-2)² + (-6)² + (-10)² = 4 + 36 + 100 = 140.
{-10,-6,-2} and (-10)² + (-6)² + (-2)² = 100 + 36 + 4 = 140.
{6,-2,-10} and (6)² + (-2)² + (-10)² = 36 + 4 + 100 = 140.
{-10,-2,6} and (-10)² + (-2)² + (6)² = 100 + 4 + 36 = 140.
{10,2,-6} and (10)² + (2)² + (-6)² = 100 + 4 + 36 = 140.
{-6,2,10} and (-6)² + (2)² + (10)² = 36 + 4 + 100 = 140.
{10,6,2} and (10)² + (6)² + (2)² = 100 + 36 + 4 = 140.
{2,6,10} and (2)² + (6)² + (10)² = 4 + 36 + 100 = 140.


There are also infinitely many irrational solutions.

If you find the missing information you can tell me what
it is in the thank-you note form below and I'll get back to 
you by email.

Edwin